OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? Answer y = 4x + 1 y = 4x - 10 y = 4x - 1 y = 4x - 5 i got A, y=4x+1

OpenStudy (chriss):

2=4(3)+b 2=12+b -10=b y=4x-10

OpenStudy (anonymous):

y-y1=m(x-x1)